OpenStudy (anonymous):

Mr Zimmer started a 6-wk exercise program. The first week he jogged 1 mile each day, the second he jogged 1 1/4 miles each day, and the third he jogged 1 1/2 mile each day. If the pattern continues, how much will he jog each day of the sixth week?

OpenStudy (anonymous):

1 * 7 + 1,25 * 7 + 1,5 * 7 + 1,75 * 7 + 2 * 7 + 2,25 * 7 = 7 + 8,75 + 10,5 + 12,25 + 14 + 15,75 = 68,25miles

Parth (parthkohli):

1 + (1/2 * 35) = 18 1/2 So answer is, 18 and a half, 19, 19 and a half, 20, 20 and a half, 21, 21 and a half

OpenStudy (anonymous):

the answers are

OpenStudy (anonymous):

1 1/4 2 1/4 6 9 3/4

Parth (parthkohli):

Oops.. I made a mistake again.. Let me do it again for you

OpenStudy (anonymous):

thanks.

Parth (parthkohli):

1 + (1/4 * 35) = 1 + 35/4 = 9 3/4

OpenStudy (anonymous):

Week n = ( 1 + 0,25 * n ) * 7 is the formula ( except for n = 1)

Parth (parthkohli):

Answer is 9 3/4

OpenStudy (anonymous):

thanks, Parth! you are like, a math genius!

OpenStudy (anonymous):

no, i don't know calculus so were ok lol

OpenStudy (anonymous):

its because you are one! ;D

OpenStudy (anonymous):

hey are you here?